OpenStudy (fernandez555):

(01.01) Given the equation 3x + 25 = 7x − 5, which order of operations completely solves for x? Add 3x, then subtract 25, lastly divide by 10 Subtract 3x, then add 5, lastly divide by 4 Add 7x, then add 25, lastly divide by 10 Subtract 7x, then add 5, lastly divide by 10

OpenStudy (the_fizicx99):

B

OpenStudy (cwrw238):

3x + 25 = 7x - 5 -3x -3x 25 = 4x - 5 + 5 + 5 30 = 4x /4 /4 30/4 = x